OpenStudy (anonymous):

Choose 3 of the polynomials from the list below and classify each one by degree and number of terms. (2 points each) 1) 9x2yz Term = 1 2) 18x2 + 5ab – 6y Term = 3 3) 8a – 5 Term = 2 4) 4x3 + 3x2 – x – 4 Term = 4 5) x + 2xyz Term = 2 6) 3x + y + z Term = 3 7) 2x3y + y2x – 3x + 4 Term = 4 8) 6x2 Term = 1 I just need help finding the Degrees, its the only thing that confuses me

OpenStudy (jackellyn):

The degree of a polynomial is the number of the highest exponent.

OpenStudy (anonymous):

Okay, well I only need 1, 5 and 7, for 1 ) I put 2 but its 1 for 5) 1 but idk the correct. for 7, i put 3 ... but idk that one either. Plz tell me is 1 correct. and the 5 and 7..

OpenStudy (jackellyn):

For polynomials in two or more variables, the degree of a term is the sum of the exponents of the variables in the term; the degree of the polynomial is again the maximum of the degrees of all terms in the polynomial.

OpenStudy (anonymous):

Ahh. Thank you.

OpenStudy (jackellyn):

1) 9x2yz Degree=4 x=2 y=1 z=1 2+1+1=4 5) x + 2xyz x x=1 -- 2xyz x=1 y=1 z=1 1+1+1=3 Degree=3 7) 2x3y + y2x – 3x + 4 2x^3y x=3 y=1 3+1=4 -- y^2x y=2 x=1 2+1=3 -- 3x x=1 Degree=4

OpenStudy (anonymous):

Thank you so much :)